The length of a rectangle is 4 times the width. The perimeter of the rectangle is 45 inches. What is the area of the rectangle. Use system of equation to solve the question. I need help finding the two equation so that i could go further and solve it. thank you.
Answer by stanbon(75887)

The length of a rectangle is 4 times the width. The perimeter of the rectangle is 45 inches. What is the area of the rectangle. Use system of equation to solve the question.
-----
width:: w
length = 4w
-----
Perimeter = 2(length + width)
45 = 2(4w+w)
2*5w = 45
2w = 9
w = 4.5 (width)
4w = 18 (length)
-----
Ans: Area = width*length = 4.5*18 = 81 sq in
-----------------
Cheers,
Stan H.
-----------------